# Learning Color Equivariant Representations

### v3

@@ D.1 Hue-shift MNIST
For this task, we compare against the Z2CNN architecture proposed in Cohen and Welling [2016]. Our hue-equivariant architecture has the same number of layers as Z2CNN, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architecture, we perform hue group pooling to yield an hue-invariant representation.
We train our equivariant networks and the conventional architectures for 5 epochs with a batch size of 128. We optimize over a cross-entropy loss using the Adam optimizer Kingma and Ba [2014] with $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We use an initial learning rate of $10^{-3}$ for the $\mathbb{Z}^{2}$ network, and $10^{-4}$ for the hue-equivariant network. We train CEConv architectures using the hyperparameters reported in Robert-Jan and van Gemert .
@@ D.2 Hue-shift 3D Shapes
For this task we compare against the Z2CNN architectures proposed in Cohen and Welling [2016]. Our hue-equivariant architectures are designed with the same network structure as the baselines, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architectures, we perform hue group pooling to yield an hue-invariant representation.
Following Wiles et al. [2021], we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section D.1.
@@ D.3 Camelyon17
For this task, we compare against the ResNet50 architecture He et al. [2016]. Our hue- and saturation-equivariant architectures have the same number of layers as ResNet50, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue- and saturation-equivariant architectures, we perform a group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 10k iterations with a batch size of 32. We optimize over a cross-entropy loss using the Adam optimizer Kingma and Ba [2014] with an initial learning rate of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train CEConv architectures using the hyperparameters reported in Robert-Jan and van Gemert .
@@ D.4 CIFAR-10
For this task, we compare against the ResNet44 architecture proposed in Cohen and Welling [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-100 Krizhevsky et al. [2009] classification dataset. The dataset consists of 60k examples, 50k of which are used for training and 10k for testing.
@@ D.5 CIFAR-100
For this task, we compare against the ResNet44 architecture proposed in Cohen and Welling [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler. We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-100 Krizhevsky et al. [2009] classification dataset. The dataset consists of 60k examples, 40k of which are used for training, 10k for validation, and 10k for testing.
@@ D.6 Caltech-101
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Caltech-101 Li et al. [2022] classification dataset. The dataset consists of 9,146 examples, $\nicefrac{{2}}{{3}}$ of which are used for training and $\nicefrac{{1}}{{3}}$ for testing.
@@ D.7 STL-101
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the STL-10 Coates et al. [2011] classification dataset. The dataset consists of 5k training examples and 8k testing examples.
@@ D.8 Stanford Cars
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Stanford Cars Krause et al. [2013] classification dataset. As the original host for the Stanford cars dataset is no longer maintained, the dataset was constructed from the images Kaggle [2018], the devkit github [2024], and the annotation files Kaggle [2023]. The dataset consists of 8,144 training examples and 8,041 testing examples.
@@ D.9 Oxford-IIT Pets
For this task, we compare against the ResNet18 architecture proposed in He et al. [2016]. Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation.
@@
We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam Kingma and Ba [2014] with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Oxford-IIT Pets  Parkhi et al. classification dataset. The dataset consists of 3,680 training examples and 3,669 testing examples.

### v4

@@ D.1 Hue-shift MNIST
We compare the classification performance of our architecture and the Z2CNN architecture proposed in Cohen & Welling (2016). Our hue-equivariant architecture has the same number of layers as Z2CNN, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architecture, we perform hue group pooling to yield an hue-invariant representation.
We train our equivariant networks and the conventional architectures for 5 epochs with a batch size of 128. We optimize over a cross-entropy loss using the Adam optimizer (Kingma & Ba, 2014) with $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We use an initial learning rate of $10^{-3}$ for the $\mathbb{Z}^{2}$ network, and $10^{-4}$ for the hue-equivariant network. We train CEConv architectures using the hyperparameters reported in (Lengyel et al., 2024).
@@ D.2 Hue-shift 3D Shapes
We compare the classification performance of our architecture and the Z2CNN architecture proposed in Cohen & Welling (2016). Our hue-equivariant architectures are designed with the same network structure as the baselines, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue-equivariant architectures, we perform hue group pooling to yield an hue-invariant representation.
Following (Wiles et al., 2022), we train the ResNet architectures for 100k iterations with a batch size of 128. We optimize over a cross-entropy loss using SGD with a learning rate of $10^{-2}$. Images are down-sampled by a factor of 2 to train the CNN architectures. CNN architectures are trained as described in Section D.1.
@@ D.3 Camelyon17
We compare the classification performance of our architecture and the ResNet50 architecture (He et al., 2016). Our hue- and saturation-equivariant architectures have the same number of layers as ResNet50, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our hue- and saturation-equivariant architectures, we perform a group pooling to yield group invariant representation.
We train our equivariant networks and the conventional architectures for 10k iterations with a batch size of 32. We optimize over a cross-entropy loss using the Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of $10^{-2}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train CEConv architectures using the hyperparameters reported in (Lengyel et al., 2024).
@@ D.4 CIFAR-10
We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-100 (Krizhevsky et al., 2009) classification dataset. The dataset consists of 60k examples, 50k of which are used for training and 10k for testing.
We compare the classification performance of our architecture and the ResNet44 architecture proposed in Cohen & Welling (2016). Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler.
@@ D.5 CIFAR-100
We evaluate the performance of our model in the presence of natural hue shifts on the CIFAR-100 (Krizhevsky et al., 2009) classification dataset. The dataset consists of 60k examples, 40k of which are used for training, 10k for validation, and 10k for testing.
We compare the classification performance of our architecture and the ResNet44 architecture proposed in Cohen & Welling (2016). Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet44, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 128. We optimize over a cross-entropy loss using SGD with an initial learning rate of $10^{-1}$ and a cosine-annealing scheduler.
@@ D.6 Caltech-101
We optimize over a cross-entropy loss using Adam (Kingma & Ba, 2014) with an initial learning rate of $10^{-2}$. We evaluate the performance of our model in the presence of natural hue shifts on the Caltech-101 (Li et al., 2022) classification dataset. The dataset consists of 9,146 examples, $\nicefrac{{2}}{{3}}$ of which are used for training and $\nicefrac{{1}}{{3}}$ for testing.
We compare the classification performance of our architecture and the ResNet18 architecture proposed in He et al. (2016). Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16.
@@ D.7 STL-101
We evaluate the performance of our model in the presence of natural hue shifts on the STL-10 (Coates et al., 2011) classification dataset. The dataset consists of 5k training examples and 8k testing examples.
We compare the classification performance of our architecture and the ResNet18 architecture proposed in He et al. (2016). Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam (Kingma & Ba, 2014) with an initial learning rate of $10^{-2}$.
@@ D.8 Stanford Cars
We evaluate the performance of our model in the presence of natural hue shifts on the Stanford Cars (Krause et al., 2013) classification dataset. As the original host for the Stanford cars dataset is no longer maintained, the dataset was constructed from the images111kaggle.com/datasets/jessicali9530/stanford-cars-dataset, the devkit222github.com/pytorch/vision/files/11644847/car_devkit.tgz, and the annotation files333kaggle.com/code/subhangaupadhaya/pytorch-stanfordcars-classification. The dataset consists of 8,144 training examples and 8,041 testing examples.
We compare the classification performance of our architecture and the ResNet18 architecture proposed in He et al. (2016). Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam (Kingma & Ba, 2014) with an initial learning rate of $10^{-2}$.
@@ D.9 Oxford-IIT Pets
We evaluate the performance of our model in the presence of natural hue shifts on the Oxford-IIT Pets (Parkhi et al., ) classification dataset. The dataset consists of 3,680 training examples and 3,669 testing examples.
We compare the classification performance of our architecture and the ResNet18 architecture proposed in He et al. (2016). Our hue-, saturation- and color-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam (Kingma & Ba, 2014) with an initial learning rate of $10^{-2}$.
D.10 Small NORB
We evaluate the performance of our model in the presence of natural luminance shifts small NORB (LeCun et al., 2004) classification dataset. The small NORB dataset consists of 48.6k 96x96 grayscale images captured under 6 different lighting conditions, 9 elevations, and 18 azimuths.
We compare the classification performance of our architecture and the ResNet18 architecture proposed in (He et al., 2016). Our luminance-equivariance architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 300 epochs, and a batch size of 16. We optimize over a cross-entropy loss using Adam (Kingma & Ba, 2014) with an initial learning rate of $10^{-2}$.
D.11 ImageNet
We evaluate the performance of our model in the presence of natural hue shifts on the ImageNet (Deng et al., 2009) classification dataset. The dataset consists 1000 categories with 14,197,122 annotated images.
We compare the classification performance of our architecture and the ResNet18 architecture proposed in (He et al., 2016). Our hue-equivariant architectures have the same number of layers as ResNet18, but with a reduced filter count to maintain a similar number of parameters at each layer. In the final layer of our networks, we perform group pooling to yield group invariant representation. We train our equivariant networks and the conventional architectures for 100 epochs, and a batch size of 64. We optimize over a cross-entropy loss using Adam (Kingma & Ba, 2014) with an initial learning rate of $10^{-3}$.
E Additional Results
E.1 Luminance shift small NORB
We demonstrate improved generalization to luminance-shifts compared to ResNet18 (He et al., 2016) on the small NORB classification dataset (LeCun et al., 2004). The small NORB dataset consist of 5 categories of objects under different lighting conditions. Variation in images reflects variation in the location and strength of the lighting. Additional details are provided in Appendix C.3.
Figure 19: Luminance equivariant feature maps. We illustrate the equivariance of our luminance-equivariant model. A luminance shift in the input image space (top-left to bottom-left), results in a feature map translation at each layer of the network (top-right to bottom-right). Corresponding feature maps are highlighted with a blue border.
Table 5: Generalization to luminance-shift. Classification error on the small NORB dataset is reported. Our model (Lum-$3$) achieves improved generalization performance over the conventional CNN model (ResNet18).
Networks A/A A/B A/C Params
ResNet18 8.32 (0.86) 37.70 (1.51) 33.88 (5.11) 11.2M
Lum-3* 7.04 (0.65) 18.45 (4.83) 25.14 (1.07) 11.1M
Luminance group and group action.
We expand the notion of color equivariance proposed in (Lengyel et al., 2024) to include equivariance to luminance shifts. In the HSL color space, luminance can be represented by a real number in the interval $[0,1]$. Using the approximations presented for construction of the discretized saturation group $S_{N}$, we similarly identify elements of the discretized luminance group $L_{N}$, with those of the integers with addition, $(\mathbb{Z},+)$.
We define the action of the luminance group on HSL images, $x\in X$ where $x:\mathbb{Z}^{2}\rightarrow\mathbb{R}^{3}$, and functions on the discrete luminance group, $y\in Y$, where $y:\mathbb{Z}^{2}\times L_{N}\rightarrow\mathbb{R}^{K}$. An element of the luminance group acts on an HSL image by the group action $\varphi_{l}:L_{N}\times X\rightarrow X$, which shifts the luminance channel of the image. Concretely, for an HSL image $x\in X$ defined as the concatenation of hue, saturation and luminance channels, i.e., $x=(x_{h},x_{s},x_{l})$, the action of an element $l_{i}$ of the luminance group $L_{N}$ is given by
$\varphi_{l}(l_{i},x)=(x_{h},x_{s},\text{min}(x_{l}+l_{i},c)),$ (34)
where $c$ is the maximum luminance value. An element of the luminance group acts on a function on the discrete luminance group by the group action $\phi_{l}:L_{N}\times Y\rightarrow Y$, which “translate” the function on the group. Concretely, for a function $f$ on the discrete luminance group $L_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $l_{i}$ in the luminance group $L_{N}$ is given by
$\phi_{l}(l_{i},f)=(f_{1+i},\dots,f_{N},\underbrace{\mathbf{0},\dots,\mathbf{0}}_{i}).$ (35)
The action of $\varphi_{l}$ on an input image, and $\phi_{l}$ on a feature map are shown in Figure 19.
Generalization to luminance-shift.
We expand the notion of color equivariance proposed in (Lengyel et al., 2024) to capture variation in luminance. The generalization performance of our luminance-equivariant models and Resnet18 (He et al., 2016) are reported in Table 5. Our luminance-equivariant model significantly outperforms the ResNet model in all out-of-distribution cases.
E.2 ImageNet
We include classification results on the ImageNet dataset (Deng et al., 2009). Additional details are provided in Appendix D.
Table 6: Generalization to color-shift. Classification error on the ImageNet dataset is reported. Our model (Hue-$3$) achieves generalization performance on-par with conventional CNN model (ResNet18) and baseline (CEConv-3).
Networks Error Params
ResNet18 30.66 11.7M
Hue-3* 30.28 11.4M
CEConv-3 30.71 11.4M
E.3 Foreground-Background Segmentation
We demonstrate the utility of our method to foreground-background segmentation on the Caltech-101 dataset (Li et al., 2022). The dataset consists of 9,146 examples with annotations containing bounding boxes and contours, which are used to generate ground truth foreground masks. The training set consists of 6941 examples, the validation set consists of 868 examples, and the test set consists of 868 examples. For uniformity of the dataset, we resize all input images and target masks to have a height and width of 224x224.
We compare the quantitative and qualitative performance of a conventional foreground-background segmentation architecture with our hue-equivariant variant. Each architecture is based on the bottleneck segmentation backbone described in Table 8 where all convolutional layers use residual connections, and each is followed by a batch normalization layer (Ioffe, 2015) and ReLU (Agarap, 2018) activation function. In the final layer, the conventional architecture uses a max pooling layer followed by a sigmoid activation function to produce the predicted foreground-background segmentation mask. Our hue-equivariant architecture is designed similarly, but uses a group pooling layer instead of a max pooling layer.
We report the quantitative performance of our hue-equivariant model and the conventional model in Table 7, and provide a qualitative comparison of the predicted foreground-background segmentation masks in Figure 20.
Table 7: Foreground-background segmentation. Segmentation error on the Caltech-101 dataset is reported. Our proposed color-equivariant convolutions achieve improved performance over the conventional convolutions.
Convolution Error Params
Conv 11.25 (0.12) 1.1M
Hue-3* 9.67 (0.23) 1.2M
Figure 20: Foreground-background segmentation. We show the foreground mask predicted by our approach and a conventional architecture. (Left to right) We show the input image, the ground truth foreground segmentation mask, the foreground segmentation mask predicted by our color-equivariant network, and the foreground segmentation mask predicted by a conventional architecture.
Table 8: Foreground-background segmentation network backbone.
Layer # Layer type Kernel size Output channels
0 input - 3
1 conv 7x7 32
2 conv 7x7 64
3 conv 7x7 128
4 conv 7x7 256
5 conv 1x1 256
6 conv 1x1 256
7 conv 1x1 256
8 deconv 7x7 128
9 deconv 7x7 64
10 deconv 7x7 32
11 deconv 7x7 3
12 pooling - 1
hello world…
hello world…